# JEL classification: C33, C38, C55

### v1

@@ 3 Theoretical Properties
Let Assumptions 17 in the Appendix hold and $\widehat{\mathbf{V}}$ be the matrix consisting of the estimated sparse eigenvectors obtained by Algorithm 2 . As $N,T\rightarrow\infty$, it holds that
$\rho(\widehat{\mathbf{V}},{\mathbf{V}}):=\|\widehat{\mathbf{V}}\widehat{\mathbf{V}}^{\prime}-{\mathbf{V}}{\mathbf{V}}^{\prime}\|_{F}=O_{p}\left(\sqrt{\frac{s^{*}\log(N)}{NT}}+\frac{s^{*}\log(N)}{NT}+\frac{1}{T}\right),$
@@
$\sqrt{\sum_{i=1}^{r}\sin^{2}(\theta_{i})}=O_{p}\left(\sqrt{\frac{s^{*}\log(N)}{NT}}+\frac{s^{*}\log(N)}{NT}+\frac{1}{T}\right),$
implying that all the principal angles will be asymptotically zero or $\pi$ if $s^{*}\log(N)=o(NT)$. On the other hand, if all directions of the estimated eigenvectors are in the same direction of the corresponding true ones.
@@
$\|\widehat{\mathbf{V}}-{\mathbf{V}}\|_{F}\asymp\rho(\widehat{\mathbf{V}},{\mathbf{V}})=O_{p}\left(\sqrt{\frac{s^{*}\log(N)}{NT}}+\frac{s^{*}\log(N)}{NT}+\frac{1}{T}\right).$
@@
$\|\widehat{\mathbf{F}}-{\mathbf{F}}\|_{F}\asymp\sqrt{T}\rho(\widehat{\mathbf{V}},{\mathbf{V}})=O_{p}\left(\sqrt{\frac{s^{*}\log(N)}{N}}+\frac{s^{*}\log(N)}{N\sqrt{T}}+\frac{1}{\sqrt{T}}\right),$
@@
(iii) When there is no sparsity, i.e., $s^{*}=T$, we have that
$\frac{1}{\sqrt{T}}\|\widehat{\mathbf{F}}-{\mathbf{F}}\|_{F}\asymp\rho(\widehat{\mathbf{V}},{\mathbf{V}})=O_{p}\left(\sqrt{\frac{\log(N)}{N}}+\frac{1}{T}\right),$
which is in line with the conventional result in factor modeling if we ignore the $\log(N)$ term. See Gao and Tsay (2024) for details.
@@
$\max_{1\leq i\leq N}\|\widehat{\boldsymbol{\lambda}}_{i}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{i}\|_{2}=O_{p}(\sqrt{\frac{s^{*}}{NT}}\log(N)+\frac{s^{*}\log(N)^{3/2}}{NT}+\sqrt{\frac{\log(N)}{T}}),$ (22)
where ${\mathbf{H}}_{s}$ is specified in (IA.18) of the Appendix. Furthermore, if $s^{*}$ is finite, then
@@
(ii) For $1\leq i\leq N$, if $s^{*}\log(N)=o(N)$, then there exists a rotation matrix ${\mathbf{H}}_{s}$ such that
@@
(ii) In practice, there are many choices for $g(N_{1},T)$ in (15). For example, we may take
@@
and
$IC(s)=\ln(R^{J}(s))+rs\frac{N_{1}+T}{N_{1}T}\ln(\frac{N_{1}T}{N_{1}+T})$ (25)
@@
(iv) When $s$ is diverging with the sample size $T$ in the sense that $s/a_{T}\asymp 1$, then $s$ should be replaced by $s/a_{T}$. For example, if $s/\sqrt{T}\asymp 1$, then the information criterion in (25) will be
$IC(s)=\ln(R^{J}(s))+r\frac{s}{\sqrt{T}}\frac{N_{1}+T}{N_{1}T}\ln(\frac{N_{1}T}{N_{1}+T}),$ (26)
which works well as shown in the simulation studies.
@@ 6 Conclusion
A Mathematical Proofs
We will use $C$ or $c$ to denote a generic constant the value of which may change at different places. We first introduce several notations used in the derivations. Let ${\mathbf{V}}={\mathbf{F}}/\sqrt{T}$, and ${\mathbf{v}}_{i}=\underline{{\mathbf{f}}}_{i}/\sqrt{T}$ be the $i$-th columns of ${\mathbf{V}}$. Define ${\mathbf{S}}_{c}={{\mathbf{F}}\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}^{\prime}{\mathbf{F}}}/{(NT)}$ and ${\mathbf{S}}_{e}={\mathbf{e}}{\mathbf{e}}^{\prime}/(NT)$. It follows from Model (3) and Assumptions 24 that
${\mathbf{S}}={\mathbf{S}}_{c}+{\mathbf{S}}_{R}=\lambda_{1}{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}+...+\lambda_{r}{\mathbf{v}}_{r}{\mathbf{v}}_{r}^{\prime}+{\mathbf{S}}_{e}+\boldsymbol{\Sigma}_{fe}+\boldsymbol{\Sigma}_{fe}^{\prime},$ (IA.1)
@@ A Mathematical Proofs
${\mathbf{B}}_{i}^{2}={\mathbf{B}}_{i}\quad\text{and}\quad{\mathbf{B}}_{i}^{\prime}{\mathbf{B}}_{i}={\mathbf{B}}_{i},$ (IA.2)
@@
Note that ${\mathbf{B}}_{0}={\mathbf{I}}_{T}$ which obviously satisfies the identities in (IA.2). From Algorithm 2, we see that ${\mathbf{B}}_{1}={\mathbf{I}}_{T}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}$ which also satisfies (IA.2) since $\widehat{\mathbf{v}}_{1}^{\prime}\widehat{\mathbf{v}}_{1}=1$. Next, we assume that ${\mathbf{B}}_{j}^{2}={\mathbf{B}}_{j}$ and ${\mathbf{B}}_{j}^{\prime}{\mathbf{B}}_{j}={\mathbf{B}}_{j}$, then, by the fact that $\widehat{\mathbf{v}}_{j}^{\prime}{\mathbf{B}}_{j}\widehat{\mathbf{v}}_{j}=1$ and an elementary argument, we can show that
@@
where we used the assumption in (IA.2) for $i=j$ and the fact that $\widehat{\mathbf{v}}_{j}^{\prime}{\mathbf{B}}_{j}\widehat{\mathbf{v}}_{j}=1$. Therefore, this proves (IA.2) for all $i$. It follows from (IA.2) that ${\mathbf{B}}_{i}$ is a symmetric matrix for all $i=1,...,r+1$. This completes the proof. $\Box$
@@
For any two unit vectors ${\mathbf{x}}$, ${\mathbf{y}}\in R^{N}$, the singular values of ${\mathbf{x}}{\mathbf{x}}^{\prime}-{\mathbf{y}}{\mathbf{y}}^{\prime}$ are
@@
where there are $N-2$ zeros.
Proof. It is a 1-dimensional case of Proposition 2.1 in Vu and Lei (2013). See also the descriptions for (3.3)–(3.4) in Gao and Tsay (2021). Specifically, note that the only non-zero singular value of ${\mathbf{x}}{\mathbf{x}}^{\prime}({\mathbf{I}}_{N}-{\mathbf{y}}{\mathbf{y}}^{\prime})$ is
$\sqrt{\mbox{tr}({\mathbf{x}}{\mathbf{x}}^{\prime}({\mathbf{I}}_{N}-{\mathbf{y}}{\mathbf{y}}^{\prime})^{2}{\mathbf{x}}{\mathbf{x}}^{\prime})}=\sqrt{\mbox{tr}\{1-({\mathbf{x}}^{\prime}{\mathbf{y}})^{2}\}}=1-({\mathbf{x}}^{\prime}{\mathbf{y}})^{2}.$
@@
where there are $N-2$ zeros. This completes the proof. $\Box$
@@
where ${\mathbf{S}}_{R}$ is defined as that in (IA.1) and $S_{1}=\{{\mathbf{v}}:\|{\mathbf{v}}\|_{2}=1\,\,\text{and}\,\,\|{\mathbf{v}}\|_{0}\leq 2s_{1}\}$.
Proof. We only focus on the case that $\widehat{\mathbf{v}}_{1}\neq{\mathbf{v}}_{1}$ as Lemmas A.4 obviously holds in such a case. For any two vectors ${\mathbf{x}}$ and ${\mathbf{y}}$, we know the trace has the property that $\mbox{tr}({\mathbf{x}}{\mathbf{y}}^{\prime})=\mbox{tr}({\mathbf{y}}^{\prime}{\mathbf{x}})$. Note that ${\mathbf{S}}_{c}{\mathbf{v}}_{1}=\lambda_{1}{\mathbf{v}}_{1}$, then
${\mathbf{S}}_{c}-\lambda_{1}{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}=({\mathbf{I}}_{n}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{S}}_{c}({\mathbf{I}}_{N}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}).$ (IA.4)
@@
$=$ $\lambda_{1}-\lambda_{1}({\mathbf{v}}^{\prime}{\mathbf{v}}_{1})^{2}-{\mathbf{v}}^{\prime}({\mathbf{I}}_{N}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{S}}_{c}({\mathbf{I}}_{N}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}},$ (IA.5)
where we use Equation (IA.4) in the last step.
Let ${\mathbf{b}}=({\mathbf{I}}_{N}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}/\|({\mathbf{I}}_{N}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}\|_{2}$, it follows that ${\mathbf{b}}^{\prime}{\mathbf{v}}_{1}=0$ and therefore, ${\mathbf{b}}\in\text{span}\{{\mathbf{v}}_{2},...,{\mathbf{v}}_{N}\}$. By (IA.1), it follows that ${\mathbf{b}}^{\prime}{\mathbf{S}}_{c}{\mathbf{b}}\leq\lambda_{2}$, which implies that
${\mathbf{v}}^{\prime}({\mathbf{I}}_{N}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{S}}_{c}({\mathbf{I}}_{N}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}\leq\lambda_{2}\|({\mathbf{I}}_{n}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}\|_{2}^{2}=\lambda_{2}-\lambda_{2}({\mathbf{v}}^{\prime}{\mathbf{v}}_{1})^{2}.$ (IA.6)
By (A) and (IA.6), we replace ${\mathbf{v}}$ by the estimator $\widehat{\mathbf{v}}_{1}$ of ${\mathbf{v}}_{1}$ and obtain
$\mbox{tr}({\mathbf{S}}_{c}({\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}))\geq(\lambda_{1}-\lambda_{2})\{1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}\}.$ (IA.7)
Now turn to the original optimization problem. Note that $\|\widehat{\mathbf{v}}_{1}\|_{2}\leq s_{1}$ and $\|{\mathbf{v}}_{1}\|_{2}\leq s_{1}$, and $\widehat{\mathbf{v}}_{1}$ is the one that maximizes the variance ${\mathbf{v}}^{\prime}{\mathbf{S}}{\mathbf{v}}$ with the constraint $\|{\mathbf{v}}\|_{2}\leq s_{1}$, then
$\mbox{tr}({\mathbf{S}}(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}))=\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{S}}\widehat{\mathbf{v}}_{1}-{\mathbf{v}}_{1}^{\prime}{\mathbf{S}}{\mathbf{v}}_{1}>0$ (IA.8)
Combining (IA.7) and (IA.8) gives
$1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}\leq\frac{1}{\lambda_{1}-\lambda_{2}}\mbox{tr}(({\mathbf{S}}-{\mathbf{S}}_{c})(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}))=\frac{1}{\lambda_{1}-\lambda_{2}}\mbox{tr}({\mathbf{S}}_{R}(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime})).$ (IA.9)
@@
It follows from (IA.9) and the above one that
$\sqrt{1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}}\leq\frac{2}{\lambda_{1}-\lambda_{2}}\|{\mathbf{D}}_{1}{\mathbf{S}}_{R}{\mathbf{D}}_{1}\|_{2}\leq\frac{2}{\lambda_{1}-\lambda_{2}}\sup_{\|{\mathbf{v}}\|_{2}=1}|{\mathbf{v}}^{\prime}{\mathbf{D}}_{1}{\mathbf{S}}_{R}{\mathbf{D}}_{1}{\mathbf{v}}|.$ (IA.10)
@@
where the set $S_{1}$ is define in Lemma A.4. Then (IA.10) reduces to
@@
$\leq$ $2\exp\left(-C\min\left(\frac{t^{2}}{N},t\right)\right),$ (IA.11)
@@
where $Z_{j}$ is defined as that in the proof of Lemma A.6. Note that $W_{j}$ and $Z_{j}$ are independent of each other and they all have finite variances. In addition, Assumption 5 implies that $Z_{j}$ are independent across $j=1,...,N$. By Assumption 7, $W_{j}Z_{j}$ satisfies condition (2.7) in Merlevède, Peligrad, and Rio (2011) with $\gamma_{2}=1$. Since $Z_{j}$ are independent over $j=1,...,N$, then condition (2.6) in Merlevède, Peligrad, and Rio (2011) is satisfied for any $\gamma_{1}>0$. By Theorem 1 of Merlevède, Peligrad, and Rio (2011), we have
$P(|\sum_{j=1}^{N}W_{j}Z_{j}|>x)\leq N\exp(-Cx^{\gamma})+\exp(-Cx^{2}/N)+\exp(-C\frac{x^{2}}{N}\exp(\frac{x^{\gamma(1-\gamma)}}{(\log(x))^{\gamma}})),$ (IA.12)
where $\gamma=(1/\gamma_{1}+1)^{-1}$. Then Lemma A.7 follows from (IA.12). $\Box$
@@
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}{\mathbf{S}}_{e}{\mathbf{v}}|\leq\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|+\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}(E{\mathbf{S}}_{e}){\mathbf{v}}|=\Pi_{1}+\Pi_{2}.$ (IA.13)
@@
$\Pi_{2}=\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}(E{\mathbf{S}}_{e}){\mathbf{v}}|\leq C/T.$ (IA.14)
Now turn to $\Pi_{1}$. Note that $S_{1}=\{{\mathbf{v}}:\|{\mathbf{v}}\|_{2}=1\,\,\text{and}\,\,\|{\mathbf{v}}\|_{0}\leq 2s_{1}\}$. Let $\mathcal{S}_{2}^{T-1}$ be the set of $N$-dimensional unit vector and $\mathcal{B}_{0}^{d}$ be the set of vectors satisfiying $\|{\mathbf{v}}\|_{0}\leq d$. Then $S_{1}=\mathcal{S}_{2}^{T-1}\cap\mathcal{B}_{0}^{2s_{1}}$.

### v2

@@ 3 Theoretical Properties
Let Assumptions 17 hold and $\widehat{\mathbf{V}}$ be the matrix consisting of the estimated sparse eigenvectors obtained by Algorithm 2 . As $N,T\rightarrow\infty$, it holds that
$\rho(\widehat{\mathbf{V}},{\mathbf{V}}):=\|\widehat{\mathbf{V}}\widehat{\mathbf{V}}^{\prime}-{\mathbf{V}}{\mathbf{V}}^{\prime}\|_{F}=O_{p}\left(\sqrt{\frac{s^{*}\log(T)}{NT}}+\frac{s^{*}\log(T)}{NT}+\frac{1}{T}\right),$
@@
$\sqrt{\sum_{i=1}^{r}\sin^{2}(\theta_{i})}=O_{p}\left(\sqrt{\frac{s^{*}\log(T)}{NT}}+\frac{s^{*}\log(T)}{NT}+\frac{1}{T}\right),$
implying that all the principal angles will be asymptotically zero or $\pi$ if $s^{*}\log(T)=o(NT)$. On the other hand, if all directions of the estimated eigenvectors are in the same direction of the corresponding true ones.
@@
$\|\widehat{\mathbf{V}}-{\mathbf{V}}\|_{F}\asymp\rho(\widehat{\mathbf{V}},{\mathbf{V}})=O_{p}\left(\sqrt{\frac{s^{*}\log(T)}{NT}}+\frac{s^{*}\log(T)}{NT}+\frac{1}{T}\right).$
@@
$\|\widehat{\mathbf{F}}-{\mathbf{F}}\|_{F}\asymp\sqrt{T}\rho(\widehat{\mathbf{V}},{\mathbf{V}})=O_{p}\left(\sqrt{\frac{s^{*}\log(T)}{N}}+\frac{s^{*}\log(T)}{N\sqrt{T}}+\frac{1}{\sqrt{T}}\right),$
@@
$\max_{1\leq i\leq N}\|\widehat{\boldsymbol{\lambda}}_{i}-{\mathbf{H}}_{s}\boldsymbol{\lambda}_{i}\|_{2}=O_{p}(\sqrt{\frac{s^{*}{{}^{2}}\log(N)}{NT^{2}}}\log(T)+\sqrt{\frac{\log(N)}{T}}),$ (22)
where ${\mathbf{H}}_{s}=(\widehat{\mathbf{F}};\widehat{\mathbf{F}})^{-1}\widehat{\mathbf{F}}^{\prime}{\mathbf{F}}$ and $s^{*}=\max(\alpha_{1},...,\alpha_{r})$. Furthermore, if $\sqrt{\frac{1}{NT^{1-2\alpha}}}\log(T)=o(1)$ where $\alpha=\max(\alpha_{1},...,\alpha_{r})$, then
@@
(ii) For $1\leq i\leq N$, if $\sqrt{\frac{1}{NT^{1-2\alpha}}}\log(T)=o(1)$, then there exists a rotation matrix ${\mathbf{H}}_{s}$ such that
@@
(ii) In practice, there are many choices for $g(N_{1},T)$ and $\alpha$ in (15). For example, we may take
@@
and $\alpha=1/2$, and
$IC(s)=\ln(R^{J}(s))+r\frac{s}{\sqrt{T}}\frac{N_{1}+T}{N_{1}T}\ln(\frac{N_{1}T}{N_{1}+T})$ (25)
@@ 6 Conclusion
Supplementary Material for “Sparse Asymptotic PCA: Identifying Sparse Latent Factors Across Time Horizon"
The Online Appendix collects the mathematical proofs of the asymptotic results in Section 3 and the descriptions of the identified sparse factors in the empirical analysis that support the main text.
Keywords: Asymptotic Principal Components, Factor Analysis, Power Method, Sparsity, High-Dimension
A Examples of Sparse Factors in Time Horizons
  1. (i)
  2. (ii)
  3. (iii)
  4. (iv)
    Monitoring and Forecasting. Stock and Watson (1998) and Forni et al. (2000b) have demonstrated that factor models provide an effective framework for monitoring economic activities, as large-dimensional economic panel data can often be summarized by a small number of factors. In this context, business cycles can be modeled through the co-movements of economic variables, which correspond to the common factors ${\mathbf{f}}_{t}$ in Model (2). When these factors exhibit sparsity, our approach allows us to track economic activities by identifying systematic risks over time. Furthermore, periods of heightened volatility in the economic panel may produce nonzero factors that are associated with specific events that influence the overall system.
B Mathematical Proofs
We will use $C$ or $c$ to denote a generic constant the value of which may change at different places. We first introduce several notations used in the derivations. Let ${\mathbf{V}}={\mathbf{F}}/\sqrt{T}$, and ${\mathbf{v}}_{i}=\underline{{\mathbf{f}}}_{i}/\sqrt{T}$ be the $i$-th columns of ${\mathbf{V}}$. Define ${\mathbf{S}}_{c}={{\mathbf{F}}\boldsymbol{\Lambda}^{\prime}\boldsymbol{\Lambda}{\mathbf{F}}^{\prime}}/{(NT)}$ and ${\mathbf{S}}_{e}={\mathbf{e}}{\mathbf{e}}^{\prime}/(NT)$. It follows from Model (3) and Assumptions 24 that
${\mathbf{S}}={\mathbf{S}}_{c}+{\mathbf{S}}_{R}=\lambda_{1}{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}+...+\lambda_{r}{\mathbf{v}}_{r}{\mathbf{v}}_{r}^{\prime}+{\mathbf{S}}_{e}+\boldsymbol{\Sigma}_{fe}+\boldsymbol{\Sigma}_{fe}^{\prime},$ (S.1)
@@ B Mathematical Proofs
${\mathbf{B}}_{i}^{2}={\mathbf{B}}_{i}\quad\text{and}\quad{\mathbf{B}}_{i}^{\prime}{\mathbf{B}}_{i}={\mathbf{B}}_{i},$ (S.2)
@@
Note that ${\mathbf{B}}_{0}={\mathbf{I}}_{T}$ which obviously satisfies the identities in (S.2). From Algorithm 2, we see that ${\mathbf{B}}_{1}={\mathbf{I}}_{T}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}$ which also satisfies (S.2) since $\widehat{\mathbf{v}}_{1}^{\prime}\widehat{\mathbf{v}}_{1}=1$. Next, we assume that ${\mathbf{B}}_{j}^{2}={\mathbf{B}}_{j}$ and ${\mathbf{B}}_{j}^{\prime}{\mathbf{B}}_{j}={\mathbf{B}}_{j}$, then, by the fact that $\widehat{\mathbf{v}}_{j}^{\prime}{\mathbf{B}}_{j}\widehat{\mathbf{v}}_{j}=1$ and an elementary argument, we can show that
@@
where we used the assumption in (S.2) for $i=j$ and the fact that $\widehat{\mathbf{v}}_{j}^{\prime}{\mathbf{B}}_{j}\widehat{\mathbf{v}}_{j}=1$. Therefore, this proves (S.2) for all $i$. It follows from (S.2) that ${\mathbf{B}}_{i}$ is a symmetric matrix for all $i=1,...,r+1$. This completes the proof. $\Box$
@@
For any two unit vectors ${\mathbf{x}}$, ${\mathbf{y}}\in R^{T}$, the singular values of ${\mathbf{x}}{\mathbf{x}}^{\prime}-{\mathbf{y}}{\mathbf{y}}^{\prime}$ are
@@
where there are $T-2$ zeros.
Proof. It is a 1-dimensional case of Proposition 2.1 in Vu and Lei (2013). See also the descriptions for (3.3)–(3.4) in Gao and Tsay (2021). Specifically, note that the only non-zero singular value of ${\mathbf{x}}{\mathbf{x}}^{\prime}({\mathbf{I}}_{T}-{\mathbf{y}}{\mathbf{y}}^{\prime})$ is
$\sqrt{\mbox{tr}({\mathbf{x}}{\mathbf{x}}^{\prime}({\mathbf{I}}_{T}-{\mathbf{y}}{\mathbf{y}}^{\prime})^{2}{\mathbf{x}}{\mathbf{x}}^{\prime})}=\sqrt{\mbox{tr}\{1-({\mathbf{x}}^{\prime}{\mathbf{y}})^{2}\}}=1-({\mathbf{x}}^{\prime}{\mathbf{y}})^{2}.$
@@
where there are $T-2$ zeros. This completes the proof. $\Box$
@@
where ${\mathbf{S}}_{R}$ is defined as that in (S.1) and $S_{1}=\{{\mathbf{v}}:\|{\mathbf{v}}\|_{2}=1\,\,\text{and}\,\,\|{\mathbf{v}}\|_{0}\leq 2s_{1}\}$.
Proof. We only focus on the case that $\widehat{\mathbf{v}}_{1}\neq{\mathbf{v}}_{1}$ as Lemmas A.4 obviously holds when $\widehat{\mathbf{v}}_{1}={\mathbf{v}}_{1}$. For any two vectors ${\mathbf{x}}$ and ${\mathbf{y}}$, we know the trace has the property that $\mbox{tr}({\mathbf{x}}{\mathbf{y}}^{\prime})=\mbox{tr}({\mathbf{y}}^{\prime}{\mathbf{x}})$. Note that ${\mathbf{S}}_{c}{\mathbf{v}}_{1}=\lambda_{1}{\mathbf{v}}_{1}$, then
${\mathbf{S}}_{c}-\lambda_{1}{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}=({\mathbf{I}}_{T}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{S}}_{c}({\mathbf{I}}_{T}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}).$ (S.4)
@@
$=$ $\lambda_{1}-\lambda_{1}({\mathbf{v}}^{\prime}{\mathbf{v}}_{1})^{2}-{\mathbf{v}}^{\prime}({\mathbf{I}}_{T}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{S}}_{c}({\mathbf{I}}_{T}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}},$ (S.5)
where we use Equation (S.4) in the last step.
Let ${\mathbf{b}}=({\mathbf{I}}_{T}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}/\|({\mathbf{I}}_{T}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}\|_{2}$, it follows that ${\mathbf{b}}^{\prime}{\mathbf{v}}_{1}=0$ and therefore, ${\mathbf{b}}\in\text{span}\{{\mathbf{v}}_{2},...,{\mathbf{v}}_{T}\}$. By (S.1), it follows that ${\mathbf{b}}^{\prime}{\mathbf{S}}_{c}{\mathbf{b}}\leq\lambda_{2}$, which implies that
${\mathbf{v}}^{\prime}({\mathbf{I}}_{T}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{S}}_{c}({\mathbf{I}}_{T}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}\leq\lambda_{2}\|({\mathbf{I}}_{T}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}){\mathbf{v}}\|_{2}^{2}=\lambda_{2}-\lambda_{2}({\mathbf{v}}^{\prime}{\mathbf{v}}_{1})^{2}.$ (S.6)
By (B) and (S.6), we replace ${\mathbf{v}}$ by the estimator $\widehat{\mathbf{v}}_{1}$ of ${\mathbf{v}}_{1}$ and obtain
$\mbox{tr}({\mathbf{S}}_{c}({\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}-\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}))\geq(\lambda_{1}-\lambda_{2})\{1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}\}.$ (S.7)
Now turn to the original optimization problem. Note that $\|\widehat{\mathbf{v}}_{1}\|_{0}\leq s_{1}$ and $\|{\mathbf{v}}_{1}\|_{0}\leq s_{1}$, and $\widehat{\mathbf{v}}_{1}$ is the one that maximizes the variance ${\mathbf{v}}^{\prime}{\mathbf{S}}{\mathbf{v}}$ with the constraint $\|{\mathbf{v}}\|_{0}\leq s_{1}$, then
$\mbox{tr}({\mathbf{S}}(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}))=\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{S}}\widehat{\mathbf{v}}_{1}-{\mathbf{v}}_{1}^{\prime}{\mathbf{S}}{\mathbf{v}}_{1}>0$ (S.8)
Combining (S.7) and (S.8) gives
$1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}\leq\frac{1}{\lambda_{1}-\lambda_{2}}\mbox{tr}(({\mathbf{S}}-{\mathbf{S}}_{c})(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime}))=\frac{1}{\lambda_{1}-\lambda_{2}}\mbox{tr}({\mathbf{S}}_{R}(\widehat{\mathbf{v}}_{1}\widehat{\mathbf{v}}_{1}^{\prime}-{\mathbf{v}}_{1}{\mathbf{v}}_{1}^{\prime})).$ (S.9)
@@
It follows from (S.9) and the above one that
$\sqrt{1-(\widehat{\mathbf{v}}_{1}^{\prime}{\mathbf{v}}_{1})^{2}}\leq\frac{2}{\lambda_{1}-\lambda_{2}}\|{\mathbf{D}}_{1}{\mathbf{S}}_{R}{\mathbf{D}}_{1}\|_{2}\leq\frac{2}{\lambda_{1}-\lambda_{2}}\sup_{\|{\mathbf{v}}\|_{2}=1}|{\mathbf{v}}^{\prime}{\mathbf{D}}_{1}{\mathbf{S}}_{R}{\mathbf{D}}_{1}{\mathbf{v}}|.$ (S.10)
@@
where the set $S_{1}$ is define in Lemma A.4. Then (S.10) reduces to
@@
$\leq$ $2\exp\left(-C\min\left(\frac{t^{2}}{N},t\right)\right),$ (S.11)
@@
where $Z_{j}$ is defined as that in the proof of Lemma A.6. Note that $W_{j}$ and $Z_{j}$ are independent of each other and they all have finite variances. In addition, Assumption 5 implies that $Z_{j}$ are independent across $j=1,...,N$. By Assumption 7, $W_{j}Z_{j}$ satisfies condition (2.7) in Merlevède et al. (2011) with $\gamma_{2}=1$. Since $Z_{j}$ are independent over $j=1,...,N$, then condition (2.6) in Merlevède et al. (2011) is satisfied for any $\gamma_{1}>0$. By Theorem 1 of Merlevède et al. (2011), we have
$P(|\sum_{j=1}^{N}W_{j}Z_{j}|>x)\leq N\exp(-Cx^{\gamma})+\exp(-Cx^{2}/N)+\exp(-C\frac{x^{2}}{N}\exp(\frac{x^{\gamma(1-\gamma)}}{(\log(x))^{\gamma}})),$ (S.12)
where $\gamma=(1/\gamma_{1}+1)^{-1}$. Then Lemma A.7 follows from (S.12). $\Box$
@@
$\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}{\mathbf{S}}_{e}{\mathbf{v}}|\leq\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}({\mathbf{S}}_{e}-E{\mathbf{S}}_{e}){\mathbf{v}}|+\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}(E{\mathbf{S}}_{e}){\mathbf{v}}|=\Pi_{1}+\Pi_{2}.$ (S.13)
@@
$\Pi_{2}=\sup_{v\in S_{1}}|{\mathbf{v}}^{\prime}(E{\mathbf{S}}_{e}){\mathbf{v}}|\leq C/T.$ (S.14)
Now turn to $\Pi_{1}$. Note that $S_{1}=\{{\mathbf{v}}:\|{\mathbf{v}}\|_{2}=1\,\,\text{and}\,\,\|{\mathbf{v}}\|_{0}\leq 2s_{1}\}$. Let $\mathcal{S}_{2}^{T-1}$ be the set of $T$-dimensional unit vector and $\mathcal{B}_{0}^{d}$ be the set of vectors satisfiying $\|{\mathbf{v}}\|_{0}\leq d$. Then $S_{1}=\mathcal{S}_{2}^{T-1}\cap\mathcal{B}_{0}^{2s_{1}}$.